An article reported that 5% of married couples in the United States are mixed racially or ethnically. Consider the population consisting of all married couples in the United States.

(a) A random sample of n = 150 couples will be selected from this population and , the proportion of couples that are mixed racially or ethnically, will be computed. What are the mean and standard deviation of the sampling distribution of ? (Round your standard deviation to four decimal places.)

mean     
standard deviation     


(b) Is it reasonable to assume that the sampling distribution of is approximately normal for random samples of size n = 150? Explain.

Yes, because  np < 10 and n(1 − p) < 10.

Yes, because  np > 10 and n(1 − p) > 10.     

No, because np < 10

.No, because np > 10.


(c) Suppose that the sample size is n = 300 rather than n = 150, as in Part (b). Does the change in sample size change the mean and standard deviation of the sampling distribution of ? What are the values for the mean and standard deviation when n = 300? (Round your standard deviation to four decimal places.)

mean     
standard deviation     


(d) Is it reasonable to assume that the sampling distribution of is approximately normal for random samples of size n = 300? Explain.

Yes, because np < 10.Yes, because np > 10.     

No, because np < 10.No, because np > 10.


(e) When n = 300, what is the probability that the proportion of couples in the sample who are racially or ethnically mixed will be greater than 0.07? (Round your answer to four decimal places.)
Answer #1

a)

here mean proportion=     μp= 0.0500
std deviation of proportion=σp=√(p*(1-p)/n)= 0.0178

b) No, because np < 10

c)

here mean proportion=     μp= 0.0500
std deviation of proportion=σp=√(p*(1-p)/n)= 0.0126

d)

Yes, because  np > 10 and n(1 − p) > 10

e)

probability =P(X>0.07)=P(Z>(0.07-0.05)/0.013)=P(Z>1.59)=1-P(Z<1.59)=1-0.9441=0.0559